# MalCVE: Malware Detection and CVE Association Using Large Language Models

### v1

@@ Abstract.
Malicious software attacks are having an increasingly significant economic impact. Commercial malware detection software can be costly, and tools that attribute malware to the specific software vulnerabilities it exploits are largely lacking. Understanding the connection between malware and the vulnerabilities it targets is crucial for analyzing past threats and proactively defending against current ones. In this study, we propose an approach that leverages large language models (LLMs) to detect binary malware, specifically within JAR files, and utilizes the capabilities of LLMs combined with retrieval-augmented generation (RAG) to identify Common Vulnerabilities and Exposures (CVEs) that malware may exploit. We developed a proof-of-concept tool called MalCVE, which integrates binary code decompilation, deobfuscation, LLM-based code summarization, semantic similarity search, and CVE classification using LLMs. We evaluated MalCVE using a benchmark dataset of 3,839 JAR executables. MalCVE achieved a mean malware detection accuracy of 97%, at a fraction of the cost of commercial solutions. It is also the first tool to associate CVEs with binary malware, achieving a recall@10 of 65%, which is comparable to studies that perform similar analyses on source code.
@@
copyright: none
@@ 1. Introduction
Existing malware detection in the industry relies on proprietary and expensive tools (e.g., VirusTotal (vendr.com, [n. d.]) and CrowdStrike (cro, [n. d.])), which often lack an explainable connection to vulnerabilities. Surveys on explainable AI in cybersecurity show that many commercial malware detectors lack transparency, making it difficult to trace alerts back to specific vulnerabilities or code patterns (The, 2024). State-of-the-art research in malware detection and vulnerability classification employs fine-tuned models to detect malware, classify its type, or identify vulnerabilities in source code snippets (Zhou et al., 2024), (Rahali and Akhloufi, 2021). However, these fine-tuning approaches do not scale to vulnerability detection in obfuscated binaries, primarily because the field lacks a large-scale dataset for this purpose. To our knowledge, no existing solution offers transparent, low-cost, and explainable malware-to-vulnerability attribution directly from obfuscated malicious binaries, especially without relying on domain-specific training.
Recent advances in LLMs and RAG offer a promising opportunity in this field (Merritt, 2025), (Lewis et al., 2021). The latest LLM models are capable of processing between 100,000 and 1,000,000 tokens as input, where one token corresponds to approximately four characters. This capability enables LLMs to analyze extensive codebases without losing contextual information (Quintero, 2024). This study investigates whether LLMs, in combination with RAG, can effectively detect malware and identify the specific vulnerabilities targeted by the malware directly from decompiled Java binaries. The research is guided by the following questions.
@@
To answer the research questions, we propose approaches that incorporate static decompilation and deobfuscation, LLM-based code summarization, semantic similarity search for relevant vulnerabilities, and LLM-based vulnerability classification. We implemented these approaches in a prototype tool called MalCVE. Its performance was evaluated based on accuracy for both RQ1 and RQ2 using the MalDICT dataset (Joyce et al., 2023), which contains 3,839 JAR executables. Furthermore, RQ2 is measured with recall@k. Results show that the tool can detect if a binary software is malicious or benign, achieving an average accuracy of 97%. The cost of detecting malware per file is low compared to commercial malware detectors, approximately 1/66 of the cost of CrowdStrike Falcon and 1/80 of the cost of ANY.RUN. The tool is also capable of listing and ranking the top 10 CVEs associated with the detected malware. The contributions of the study are:
@@
  • Most components used to implement the approaches in the proof-of-concept tool MalCVE are open-source, with the exception of the LLMs, which may be commercially available models. This design lowers the barrier to entry for individual researchers, academic institutions, and small enterprises to perform malware analysis, not only by reducing costs but also by minimizing the need for specialized expertise.
The remainder of this paper is organized as follows. Section 2 introduces related work. Section 3 presents the proposed approach and describes the implementation of MalCVE. Section 4 details the evaluation design and presents the results. Section 5 discusses the results. Section 6 concludes the study and outlines directions for future research.
@@ 2.1. Malware Detection Using LLMs
Rahali and Akhloufi (2021) proposed MalBERT, a malware detection and categorisation solution for Android APKs. Their approach fine-tunes a pre-trained model on 12,000 benign applications and 10,000 malicious applications. The final solution achieved 97.6% for binary classification and 91% for malware type classification.
Quintero (2024) presents an in-depth test of Gemini 1.5 Pro, highlighting its application in malware detection. The study also presents analysis examples that showcase the model’s proficiency in identifying malicious code and summarizing its behavior and key functionality. The model successfully detected a zero-day attack that had not been previously flagged by antivirus software. This demonstrates that LLMs have capabilities that surpass those of traditional methods, such as antivirus software. Quintero et al. (2024) provides an example demonstrating the LLM’s capability to identify false positives. A file flagged by an antivirus vendor on VirusTotal (vendr.com, [n. d.]) was correctly identified as benign code intended for browser updates after a 1.51-second analysis by the LLM. A key limitation of the paper was that it presented only a few examples without systematic quantitative evaluations.
Li et al. (Xiang Li, 2023) developed a novel ransomware detection methodology utilizing open-source LLMs and image analysis of portable executable (PE) files. Their research focused on four types of ransomware and four types of benignware. The results demonstrate the capability of LLMs to distinguish malware from benignware, with false positive and false negative rates below 7% across the four types analyzed. Although the results of this research are promising, the study consists of only 221 files across the eight file types. The dataset is small and imbalanced, which could skew the results. Walton et al. (2025) used OpenAI’s GPT-4o-mini model to generate summaries for code at the function, class, and package levels for Android APK malware. Using these summaries, the model attempts to determine whether the Android APKs are malicious or benign. Using prompt engineering alone, their research achieved 77% classification accuracy. However, the dataset consists of only 200 files. Additionally, the research does not address code obfuscation techniques (Zhang et al., 2023). Code obfuscation is the process of transforming code into a functionally equivalent but difficult-to-understand version, which is a common approach for attackers to confuse malware detectors and can result in malicious files being classified as benign.
@@ 2.2. Vulnerability Detection in Source Code Using LLM and RAG
Du et al. (2024) found that current machine learning-based solutions struggle to achieve a deep understanding of vulnerability semantics, particularly when distinguishing between subtly different but semantically distinct code pairs. Their solution addresses this problem by integrating RAG, which serves as an external knowledge source. A tool named Vul-RAG was developed to demonstrate the idea, achieving a 61% accuracy on the benchmark dataset they created. Furthermore, their user study showed that LLM-generated knowledge increased manual vulnerability detection by 17%, confirming its value. Although Vul-RAG outperformed state-of-the-art solutions, false negatives often happened because of missing or inaccurate vulnerability knowledge, i.e., there was either insufficient or wrong context for the LLM.
Cao and Jun (2024) introduced LLM-CloudSec, an unsupervised framework that employs LLMs for fine-grained vulnerability detection and analysis in cloud applications. Their aim is to overcome the limitations of traditional static and dynamic analysis and existing machine learning-based methods by incorporating RAG and Common Weakness Enumeration (CWE) as an external knowledge base. LLM-CloudSec comprises two collaborating LLM agents: a detection agent and an analysis agent. The detection agent uses few-shot learning with embedded CWE category names and numbers to identify whether a given code snippet contains a vulnerability and to classify it into one or more CWE categories via RAG. Upon positive detection, the analysis agent retrieves a detailed description of the CWE, performs line-level localization, and generates an in-depth natural language analysis of the vulnerability. Both agents rely on GPT-4-0613 as the underlying LLM. LLM-CloudSec was tested on a dataset consisting of 10 instances of vulnerable code for 60 distinct CWEs of the Juliet C++ dataset. The detection agent achieved 70% accuracy in classifying CWE vulnerabilities.
@@ 3. Our approach and MalCVE implementation
Although existing studies have explored the potential of using LLMs to identify binary malware, they are limited to evaluations on small or imbalanced datasets and have not investigated the impact of code obfuscation. In terms of associating malware with vulnerabilities, existing studies are limited to source code analysis and CWE. While identifying CWEs is valuable for addressing the root causes of vulnerabilities, identifying CVEs is more valuable, as it provides actionable insights for remediation. Our approach aims to address these limitations. This section outlines the design principles of our approach and describes the implementation of the proof-of-concept tool, MalCVE.
@@ 3.1.1. Decompilation
The first step in our process is decompiling the JAR executable, which is a prerequisite for the subsequent summarization step. The most widely adopted commercial disassembler and decompiler is IDA Pro [^1] https://hex-rays.com/ida-pro, which supports a wide array of architectures and file formats. IDA Pro is widely regarded as the industry standard for software reverse engineering (Kaspersky, 2025). It employs signature matching FLIRT [^2] https://docs.hex-rays.com/user-guide/signatures/flirt, which recognizes standard library functions generated by supported compilers (FLI, 2025) and Lumina [^3] https://hex-rays.com/lumina, which tracks metadata about some widely recognizable functions, such as their names, prototypes, or operand types (lum, [n. d.]). Binary Ninja [^4] https://binary.ninja/ provides a simplified GUI and a well-defined intermediate language, facilitating both manual and automated analysis. Its built-in decompiler and multi-language SDK (Python, C++, Rust) make it particularly popular among vulnerability researchers seeking to script complex analyses. While IDA Pro is considered state-of-the-art, several specialized open-source alternatives are available for specific programming languages. CFR [^5] https://www.benf.org/other/cfr/ is an open-source decompiler specialized for Java. CFR supports the decompilation of modern Java features, yet it is implemented in Java 6, allowing compatibility with older Java files. In addition, CFR is able to turn other JVM languages, such as Kotlin or Scala, back into Java (Benfield, [n. d.]). According to Strobel (Strobel, [n. d.]), CFR is a better open-source decompiler for obfuscated JARs. Procyon [^6] https://github.com/ststeiger/procyon is another open-source decompiler for Java (Strobel, [n. d.]). This decompiler also excels at newer Java features. Ghidra [^7] https://ghidra-sre.org/, released by the NSA in 2019, has rapidly gained popularity in the open-source community. Its decompiler supports a wide range of architectures and translates binaries into human-readable pseudo C/C++ code.
We employ a combination of two decompilers—CFR 0.152 and Procyon 0.6.0—to leverage open-source solutions and minimize operational costs. Using two distinct decompilers provides redundancy, ensuring that if one fails to decompile a given JAR file, the other may succeed. Ghidra was also tested for this use case, as it is a popular open-source choice. However, CFR and Procyon are more suitable for JAR files than Ghidra, as they generate actual Java source code rather than pseudocode representations. The system attempts to use the CFR decompiler as the primary option. If CFR fails to decompile the JAR, the system falls back to Procyon. In most cases, at least one of the decompilers successfully processes the file. If both fail, the file is excluded from further analysis. The decompiled code is temporarily stored until the analysis is complete. Subsequently, it is removed from the system to reduce storage overhead.
@@ 3.1.2. Deobfuscation
In the second step, a deobfuscation tool was applied. Deobfuscating the code ensures that the LLM can sufficiently understand the program’s behaviour, generate accurate search queries based on used libraries, methods, and classes, and determine the nature of the code.
Java deobfuscation tools aim to reverse common obfuscation techniques used to hinder reverse engineering. Several general-purpose tools are available for this purpose, such as Java Deobfuscator [^8] https://github.com/java-deobfuscator/deobfuscator. Many of these deobfuscators have limitations, such as outdated codebases, lack of maintenance, limited support for specific obfuscators, or poor scalability. For instance, the Java Deobfuscator repository has not received updates in over two years. For this task, Java Deobfuscator was tested because it is open source. However, testing revealed that Java Deobfuscator failed to detect obfuscation patterns in 30 malicious JAR files from the MalDICT dataset (Joyce et al., 2023). Due to the limited availability of open-source deobfuscators for Java, we created a custom deobfuscation tool using Java and the JavaParser library (javaparser.org, [n. d.]). Our deobfuscator specifically looks for simple string obfuscation techniques that hide string literals and attempts to resolve those static strings. This approach does not address all obfuscation techniques and may not succeed in every case. However, it allows us to assess its impact on LLMs’ ability to understand the code by potentially revealing obfuscated strings. Figure 1 highlights the differences in the code before and after applying our deobfuscator. Red lines indicate the original code (decompiled but not yet deobfuscated), while green lines represent the code after applying deobfuscation.
@@ 3.1.3. Summarization of the code
The third step is to summarize the code using a LLM. This step answers RQ1 by including a verdict (i.e., classifying the file as Benign, Suspicious, or Malicious, and providing a confidence score and rationale) in the summary. This step also contributes to answering RQ2 by extracting knowledge and keywords that can be searched for in the vector database, thus extracting relevant CVEs. Summarizing the code is done by prompting the LLM with the decompiled and deobfuscated code, along with an instruction prompt. Figure 2 presents an example input prompt. This input enables the extraction of technical details that can be matched with CVE descriptions. The LLM generates a structured response, defined by a JSON schema containing specific properties. The response includes a verdict, summary, observed activities, indicators of compromise (IoCs), utilized libraries, and relevant CVE search queries. An example output is shown in Figure 3.

### v2

@@ Abstract.
Malicious software attacks are having an increasingly significant economic impact. Commercial malware detection software can be costly, and tools that attribute malware to the specific software vulnerabilities it exploits are largely lacking. Understanding the connection between malware and the vulnerabilities it targets is crucial for analyzing past threats and proactively defending against current ones. In this study, we propose an approach that leverages large language models (LLMs) to detect binary malware, specifically within JAR files, and uses LLM capabilities combined with retrieval-augmented generation (RAG) to identify Common Vulnerabilities and Exposures (CVEs) that malware may exploit. We developed a proof-of-concept tool, MalCVE, that integrates binary code decompilation, deobfuscation, LLM-based code summarization, semantic similarity search, and LLM-based CVE classification. We evaluated MalCVE using a benchmark dataset of 3,839 JAR executables. MalCVE achieved a mean malware-detection accuracy of 97%, at a fraction of the cost of commercial solutions. In particular, the results demonstrate that LLM‑based code summarization enables highly accurate and explainable malware identification. MalCVE is also the first tool to associate CVEs with binary malware, achieving a recall@10 of 65%, which is comparable to studies that perform similar analyses on source code.
@@
$c$s: Security and privacy Software security engineering
@@ 1. Introduction
Existing malware detection in the industry relies on proprietary and expensive tools (e.g., VirusTotal (vendr.com, ) and CrowdStrike (5)), which often lack an explainable connection to vulnerabilities. Surveys on explainable AI in cybersecurity show that many commercial malware detectors lack transparency, making it difficult to trace alerts back to specific vulnerabilities or code patterns (23). State-of-the-art research in malware detection and vulnerability classification employs fine-tuned models to detect malware, classify its type, or identify vulnerabilities in source code snippets (Zhou et al., 2024), (Rahali and Akhloufi, 2021). However, these fine-tuning approaches do not scale to vulnerability detection in obfuscated binaries, primarily because the field lacks a large-scale dataset for this purpose. To our knowledge, no existing solution offers transparent, low-cost, and explainable malware-to-vulnerability attribution directly from obfuscated malicious binaries, especially without relying on domain-specific training.
Recent advances in LLMs and RAG offer a promising opportunity in this field (Merritt, 2025), (Lewis et al., 2021). The latest LLM models can process between 100,000 and 1,000,000 tokens, where each token corresponds to approximately 4 characters. This capability enables LLMs to analyze extensive codebases without losing contextual information (Quintero, 2024). This study investigates whether LLMs, in combination with RAG, can effectively detect malware and identify the specific vulnerabilities it targets directly from decompiled Java binaries. The research is guided by the following questions.
@@
To answer the research questions, we propose approaches that incorporate static decompilation and deobfuscation, LLM-based code summarization, semantic similarity search for relevant vulnerabilities, and LLM-based vulnerability classification. We implemented these approaches in a prototype tool called MalCVE. Its performance was evaluated using accuracy metrics for both RQ1 and RQ2 on the MalDICT dataset (Joyce et al., 2023), which contains 3,839 JAR executables. Furthermore, RQ2 is measured with recall@k. Results show that MalCVE can detect if a binary software is malicious or benign, achieving an average accuracy of 97%. The cost of detecting malware per file is low compared to commercial malware detectors, approximately 1/66 of the cost of CrowdStrike Falcon  [^1] https://www.crowdstrike.com/en-us/ and 1/80 of the cost of ANY.RUN  [^2] https://any.run/. The tool can also list and rank the top 10 CVEs associated with the detected malware. The contributions of this study are:
@@
  • Most components used to implement the approaches in the proof-of-concept tool MalCVE are open source, except for the LLMs, which may be commercial models. This design lowers the barrier to entry for individual researchers, academic institutions, and small enterprises to perform malware analysis, not only by reducing costs but also by minimizing the need for specialized expertise.
The remainder of this paper is organized as follows. Section 2 introduces related work. Section 3 presents the proposed approach. Section 4 details the evaluation design and presents the results. Section 5 discusses the results. Section 6 concludes the study and outlines future research.
@@ 2.1. Malware Detection Using LLMs
Rahali and Akhloufi (2021) proposed MalBERT, a malware detection and categorisation solution for Android APKs. Their approach fine-tunes a pre-trained model on 12,000 benign applications and 10,000 malicious applications. The final solution achieved 97.6% accuracy for binary classification and 91% for malware type classification.
Quintero (2024) presents an in-depth test of Gemini 1.5 Pro, highlighting its application in malware detection. The study also presents examples that showcase the model’s proficiency in identifying malicious code and summarizing its behavior and key functionality. The model successfully detected a zero-day attack that antivirus software had not previously flagged. This demonstrates that LLMs have capabilities that surpass those of traditional methods, such as antivirus software. Quintero et al. (2024) provides an example demonstrating LLM’s capability to identify false positives. A file flagged by an antivirus vendor on VirusTotal (vendr.com, ) was correctly identified as benign by the LLM. A key limitation of the paper (Quintero et al., 2024) was that it presented only a few examples without systematic quantitative evaluations.
Li et al. (Xiang Li, 2023) developed a novel ransomware-detection methodology that uses open-source LLMs and image analysis of portable executable (PE) files. Their research focused on four types of ransomware and four types of benignware. The results demonstrate the capability of LLMs to distinguish malware from benignware, with false-positive and false-negative rates below 7% across the four types analyzed. Although the results of this research are promising, the study comprises only 221 files across eight file types. The dataset is small and imbalanced, which could skew the results. Walton et al. (2025) used OpenAI’s GPT-4o-mini model to generate summaries for code at the function, class, and package levels for Android APK malware. Using these summaries, the model attempts to determine whether the Android APKs are malicious or benign. Using prompt engineering alone, their research achieved 77% classification accuracy. However, the dataset consists of only 200 files. Additionally, the research does not address code obfuscation techniques (Zhang et al., 2023). Code obfuscation is the process of transforming code into a functionally equivalent but difficult-to-understand form, a common approach for attackers to confuse malware detectors and cause malicious files to be classified as benign.
@@ 2.2. Vulnerability Detection in Source Code Using LLM and RAG
Du et al. (2024) found that current machine learning-based solutions struggle to achieve a deep understanding of vulnerability semantics, particularly when distinguishing between subtly different but semantically distinct code pairs. Their solution addresses this problem by integrating RAG. A tool named Vul-RAG was developed to demonstrate the idea, achieving a 61% accuracy on the benchmark dataset they created. Furthermore, their user study showed that LLM-generated knowledge increased manual vulnerability detection by 17%, confirming its value. Although Vul-RAG outperformed state-of-the-art solutions, false negatives often occurred due to missing or inaccurate vulnerability knowledge, i.e., insufficient or incorrect context for the LLM.
Cao and Jun (2024) introduced LLM-CloudSec, an unsupervised framework that employs LLMs for fine-grained vulnerability detection and analysis in cloud applications. Their aim is to overcome the limitations of traditional static and dynamic analysis and of existing machine learning-based methods by incorporating RAG and the Common Weakness Enumeration (CWE) as external knowledge bases. LLM-CloudSec comprises two collaborating LLM agents: a detection agent and an analysis agent. The detection agent uses few-shot learning with embedded CWE category names and numbers to identify whether a given code snippet contains a vulnerability and to classify it into one or more CWE categories via RAG. Upon a positive detection, the analysis agent retrieves a detailed description of the CWE, performs line-level localization, and generates an in-depth natural-language analysis of the vulnerability. Both agents rely on GPT-4-0613 as the underlying LLM. LLM-CloudSec was tested on a dataset comprising 10 instances of vulnerable code across 60 distinct CWEs from the Juliet C++ dataset. The detection agent achieved 70% accuracy in classifying CWE vulnerabilities.
@@ 3. Our approach and MalCVE implementation
Although existing studies have explored the potential of LLMs for identifying binary malware, they are limited to evaluations on small or imbalanced datasets and have not investigated the impact of code obfuscation. In terms of associating malware with vulnerabilities, existing studies are limited to source code analysis and CWE. While identifying CWEs is valuable for addressing the root causes of vulnerabilities, identifying CVEs is more valuable, as it provides actionable insights for remediation. Our approach aims to address these limitations. This section outlines the design principles of our approach and describes the implementation of the proof-of-concept tool, MalCVE.
@@ 3.1.1. Decompilation
The first step in our process is decompiling the JAR executable, which is a prerequisite for the subsequent summarization step. The most widely adopted commercial disassembler and decompiler is IDA Pro  [^3] https://hex-rays.com/ida-pro, which supports a wide array of architectures and file formats. IDA Pro is widely regarded as the industry standard for software reverse engineering (Kaspersky, 2025). It employs signature matching FLIRT  [^4] https://docs.hex-rays.com/user-guide/signatures/flirt, which recognizes standard library functions generated by supported compilers (6) and Lumina  [^5] https://hex-rays.com/lumina, which tracks metadata about some widely recognizable functions, such as their names, prototypes, or operand types (13). Binary Ninja  [^6] https://binary.ninja/ provides a simplified GUI and a well-defined intermediate language, facilitating both manual and automated analysis. Its built-in decompiler and multi-language SDK (Python, C++, Rust) make it particularly popular among vulnerability researchers seeking to script complex analyses. While IDA Pro is considered state-of-the-art, several specialized open-source alternatives are available for specific programming languages. CFR  [^7] https://www.benf.org/other/cfr/ is an open-source decompiler specialized for Java. CFR supports the decompilation of modern Java features, yet it is implemented in Java 6, allowing compatibility with older Java files. In addition, CFR can convert other JVM languages, such as Kotlin or Scala, back into Java (Benfield, ). According to Strobel (Strobel, ), CFR is a better open-source decompiler for obfuscated JARs. Procyon  [^8] https://github.com/ststeiger/procyon is another open-source decompiler for Java (Strobel, ). This decompiler also excels at newer Java features. Ghidra  [^9] https://ghidra-sre.org/, released by the NSA in 2019, has rapidly gained popularity in the open-source community. Its decompiler supports a wide range of architectures and translates binaries into human-readable pseudo C/C++ code.
We employ a combination of two decompilers—CFR 0.152 and Procyon 0.6.0—to leverage open-source solutions and minimize operational costs. Using two distinct decompilers provides redundancy, ensuring that if one fails to decompile a given JAR file, the other may succeed. Ghidra was also tested for this use case, as it is a popular open-source choice. However, CFR and Procyon are better suited to JAR files than Ghidra, as they generate actual Java source code rather than pseudocode. MalCVE attempts to use the CFR decompiler as the primary option. If CFR fails to decompile the JAR, the system falls back to Procyon. In most cases, at least one of the decompilers successfully processes the file. If both fail, the file is excluded from further analysis.
@@ 3.1.2. Deobfuscation
In the second step, a deobfuscation tool was applied. Deobfuscating the code ensures that the LLM can sufficiently understand the program’s behavior, generate accurate search queries based on the libraries, methods, and classes used, and determine the nature of the code.
Java deobfuscation tools aim to reverse common obfuscation techniques used to hinder reverse engineering. Several general-purpose tools are available for this purpose, such as Java Deobfuscator  [^10] https://github.com/java-deobfuscator/deobfuscator. Many of these deobfuscators have limitations, such as outdated codebases, lack of maintenance, limited support for specific obfuscators, or poor scalability. For instance, the Java Deobfuscator repository has not received updates in over two years. For this task, Java Deobfuscator was tested because it is open source. However, testing revealed that Java Deobfuscator failed to detect obfuscation patterns in 30 malicious JAR files from the MalDICT dataset (Joyce et al., 2023). Due to the limited availability of open-source deobfuscators for Java, we created a custom deobfuscation tool using Java and the JavaParser library (javaparser.org, ). Our deobfuscator parses Java source code, walks its AST (Abstract Syntax Tree), evaluates constant-like expressions (especially string-related ones), and replaces obfuscated expressions in the source with their computed literal values. The approach effectively deobfuscates Java code by resolving and inlining constant string expressions. Although the approach does not address all obfuscation techniques and may not succeed in every case, it allows us to assess LLMs’ ability to understand the code by potentially revealing obfuscated strings. Figure 1 provides an example and highlights the differences in the code before and after applying our deobfuscator. Red lines indicate the original code (decompiled but not yet deobfuscated), while green lines represent the code after applying deobfuscation.
@@ 3.1.3. Summarization of the code
The third step is to summarize the code using an LLM. This step answers RQ1 by including a verdict (i.e., classifying the file as Benign or Malicious, and providing a confidence score and rationale) in the summary. This step also contributes to answering RQ2 by extracting knowledge and keywords that can be searched in the vector database, thereby identifying relevant CVEs. Summarizing the code is done by prompting the LLM with the decompiled and deobfuscated code, along with an instruction prompt. Figure 2 presents an example input prompt. This input enables the extraction of technical details that can be matched with CVE descriptions. The LLM generates a structured response, defined by a JSON schema containing specific properties. The response includes a verdict, summary, observed activities, indicators of compromise (IoCs), utilized libraries, and relevant CVE search queries. An example output is shown in Figure 3.